OpenStudy (anonymous):

The pair of points (6, y) and (10, –1) are on a line with slope of 1 over 4. What is the value of y?

OpenStudy (anonymous):

-2

OpenStudy (across):

\[\frac{-1-y}{10-6}=\frac{1}{4}\]solve for y

OpenStudy (karatechopper):

y+1/6-10=1/4 y+1/-4=1/4 -4(y+1/4)=(1/4)-4 y+1=-1 y+1-1=-1-1 y=-2